MICROWAVE SPECTRA AND GEOMETRIES OF H_{2}C_{2}⋯AgCl AND H_{2}C_{2}⋯CuCl

\underline{N. R. WALKER}, S. L. STEPHENS, W. MIZUKAMI, D. P. TEW AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

Pure rotational spectra of the vibrational ground states of H_{2}C_{2}⋯AgCl and H_{2}C_{2}⋯CuCl have been measured by chirpedpulse FTMW spectroscopy. Each complex is generated via laser ablation of the metal in the presence of small percentages of CCl_{4} and C_{2}H_{2} in argon. The complexes are stabilized and interrogated in the cold environment of a supersonic jet. Rotational constants (B_{0}, C_{0}) and the centrifugal distortion constant, Δ_{J} , have been measured for six isotopologues of H_{2}C_{2}⋯AgCl and three isotopologues of H_{2}C_{2}⋯CuCl with substitutions at the metal, chlorine and carbon atoms in each case. The spectrum of each complex is consistent with a C_{2v} structure in which the metal atom is coordinated by the π-orbital of ethyne. The measured rotational constants allow determination of the length of the bond between the metal and chlorine atoms, r(M—Cl), and the distance between the metal atom and the centre of the ethyne double bond, r(M—*). Nuclear quadrupole coupling constants have been measured for the chlorine atom in each complex and also for copper in H_{2}C_{2}⋯CuCl.